# Accuracy of Charge Densities in Electronic Structure Calculations

### v2

@@ I Introduction
In recent years, machine learning potentials have become an important addition to ab initio methods, reproducing electronic structure calculations with high accuracy at only a fraction of the computational cost [4, 5, 6, 7, 8, 9, 10, 11]. Many recent machine learning potentials [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] include effects like long range charge transfer, which necessitates the ability to accurately model charge densities using ab initio methods. Due to the high computational costs it is generally not feasible to compute reference data such as atomic Hirshfeld charges [23] using highly accurate methods like coupled cluster with single and double excitations (CCSD) [24, 25]. Instead, more approximate methods like DFT [2, 3] are typically used, trading some accuracy for increased computational efficiency. However, the accuracy of charge densities is strongly influenced by the choice of the exchange correlation functional and even the most accurate DFT functionals suffer from an electron self-interaction error [26, 27, 28, 29] that impacts the accuracy of the charge density. Medvedev et al. [30] even claim that after the early 2000s the density error in DFT functionals is in fact increasing. Controlling the self-interaction error in approximate forms of the exchange correlation functional is a highly active field of research, with many approaches aiming to remove this source of error  [31, 32, 33, 34, 35].
In this paper, we benchmark widely distributed Gaussian basis sets in combination with various DFT functionals and compare the accuracy of the obtained charge densities with CCSD calculations using PySCF [36]. Specifically, we assess the performance of different basis sets, including polarization consistent [37] and correlation consistent[38] sets, and analyze the resulting densities in terms of the numerical stability of Hirshfeld charge partitioning.
@@ II.1 Hirshfeld charges
The Hirshfeld charge [23] $q_{i}$ of atom $i$ is defined as
@@ II.2 Gaussian basis sets
The biggest advantage of Gaussian orbitals is that overlap, kinetic, nuclear repulsion and coulomb integrals can be solved analytically.
Disadvantages of Gaussian basis sets are that a large number of Gaussian functions is required to accurately represent the nuclear cusps in wave functions and Gaussian basis sets are not orthonormal meaning that convergence of the basis set to the complete basis set limit can not necessarily be achieved by adding more basis functions. Finally, the condition number of the overlap matrix increases with the addition of Gaussian basis functions, leading to numerical problems that often limit the number of basis functions that can be used in a calculation.
@@ II.3 Multi resolution Wavelets
The Multi wavelet approach can be coupled by a separable form to represent Greens functions of the Poisson and Helmholtz equation [47] which enables efficient Hartree Fock [46] and DFT [45] calculations.
These features are implemented in the software package MRChem [41, 42] and this code will be used for all wavelet based calculations in this paper.
@@ III.1 Basis set errors of charge densities
First, the error of the charge density for polarization consistent Gaussian basis sets [37] that are optimized for DFT calculations is analyzed. In Fig. 1, the difference of $\sigma$ between an basis-set-error-free multi wavelet charge density and a charge density obtained with polarization consistent basis sets [37] is shown. The charge density was calculated with the PBE functional for a single neon atom. The “aug” prefix indicates the addition of diffuse functions in the Gaussian basis set. The aug-PC3 and aug-PC4 basis sets agree well with the multi wavelet reference calculation. For comparison, a charge density that was obtained using a correlation consistent basis set (aug-cc-pV5Z) is added in Fig. 1 which gives the same level of accuracy in the charge density. The aug-PC3 contains fewer basis functions than the aug-cc-pV5z basis set, the aug-PC4 more. In Fig. 2 the basis set error of charge densities obtained with a correlation consistent basis [38] sets is shown. Also the aug-cc-pV5Z basis set matches the multi wavelet result almost perfectly. It is therefore recommended to use this basis set whenever highly accurate charge densities are necessary with wave function based methods. The aug-PC3 and aug-PC4 offer the same level with a similar number of basis functions, the aug-PC3 is even a bit smaller than the aug-cc-pV5Z and is therefore an interesting alternative to the correlation consistent basis set when DFT is used. In Fig. 3 it is shown that the aug-cc-pV5Z basis set also produces highly accurate charge densities for coupled cluster calculations with single and double excitations. The error of charge densities with the cc-pVNZ and the PC basis is analyzed in Fig. 4 for a methanol molecule. The reference charge density was obtained using a multi wavelet DFT calculation with the PBE functional. As a measure of error, in this case the absolute value of the difference in Hirshfeld charges is used. Here, the addition of diffuse Gaussian basis functions greatly improves the accuracy and once again, the aug-cc-pV5Z basis is able to produce highly accurate charge densities. The aug-PC3 and aug-PC4 basis sets also produce highly accurate Hirshfeld charges, the aug-PC4 appears to be the most accurate basis set in this comparison. However, the calculation using the aug-PC4 basis set was difficult to converge since the condition number of the overlap matrix from the basis functions was larger than $10^{8}$. aug-PC3, aug-PC4 and the aug-cc-pV5Z offer sufficient precision to accurately compare charge densities obtained with different functionals. Given that DFT results are being compared to CCSD calculations, it is more consistent to use the correlation-consistent aug-cc-pV5Z basis set, which is specifically designed for post-Hartree–Fock methods. Therefore, in the remainder of this paper, the aug-cc-pV5Z basis set will be used for comparing various exchange-correlation functionals.
@@ III.2 Accuracy measures of charge densities
The integrals required to calculate the norms were evaluated numerically with the integration grids of PySCF using the accurate grid level. The reference charge densities required for the calculation of the Hirshfeld charges were taken from Abinit [49]. In Fig. 6, these measures are shown for all functionals. The functionals tested are:

### v3

@@ I Introduction
In recent years, machine learning potentials have become an important addition to ab initio methods, reproducing electronic structure calculations with high accuracy at only a fraction of the computational cost [4, 5, 6, 7, 8, 9, 10, 11]. Many recent machine learning potentials [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] include effects like electrostatics and long range charge transfer, which necessitates the ability to accurately model charge densities using ab initio methods. Notably, some approaches [23, 24, 25, 26] focus on directly modeling the electron density itself, using it for constructing transferable and interpretable potentials. These advancements highlight the critical role of accurate charge densities. Due to the high computational costs it is generally not feasible to compute reference data such as atomic Hirshfeld charges [27] using highly accurate methods like coupled cluster with single and double excitations (CCSD) [28, 29]. Instead, more approximate methods like DFT [2, 3] are typically used, trading some accuracy for increased computational efficiency. However, the accuracy of charge densities is strongly influenced by the choice of the exchange correlation functional and even the most accurate DFT functionals suffer from an electron self-interaction error [30, 31, 32, 33] that impacts the accuracy of the charge density. Medvedev et al. [34] even claim that after the early 2000s the density error in DFT functionals is in fact increasing. Additionally, several studies [35, 36, 37, 38] have analyzed atomic charge densities to identify and understand the error sources in DFT, focusing on the impact of functionals and electron self interaction errors. Controlling the self interaction error in approximate forms of the exchange correlation functional is a highly active field of research, with many approaches aiming to remove this source of error [39, 40, 41, 42, 43].
In this paper, we benchmark widely distributed Gaussian basis sets in combination with various DFT functionals and compare the accuracy of the obtained charge densities of molecules with CCSD calculations using the PySCF code [44]. CCSD with pertubative triple excitations is not used here because the pertubative energy correction does not affect the charge density. Specifically, we assess the performance of different basis sets, including polarization consistent [45] and correlation consistent[46] sets, and analyze the resulting densities in terms of the numerical stability of Hirshfeld charge partitioning. Such comparisons are essential for evaluating the reliability of computational approaches and guiding the selection of appropriate methodologies [47, 48, 49, 50].
@@ II.1 Hirshfeld charges
Assigning partial charges to atoms in molecules is an important task in computational chemistry, as these charges serve several purposes: they can be used as parameters in force fields and machine learned potentials, or to qualitatively analyze the electronic structure of the system under study. Hirshfeld charges are one of many methods [51, 52, 53, 54, 55, 56, 57] to partition the electron density into atomic partial charges. The Hirshfeld charge [27] $q_{i}$ of atom $i$ is defined as
@@ II.2 Gaussian basis sets
The biggest advantage of Gaussian orbitals is that overlap, kinetic, nuclear repulsion and Coulomb integrals can be solved analytically. Disadvantages of Gaussian basis sets are that a large number of Gaussian functions is required to accurately represent the nuclear cusps in wave functions and Gaussian basis sets are not orthonormal meaning that convergence of the basis set to the complete basis set limit can not necessarily be achieved by adding more basis functions. Finally, the condition number of the overlap matrix increases with the addition of Gaussian basis functions, leading to numerical problems that often limit the number of basis functions that can be used in a calculation.
@@ II.3 Multi resolution wavelets
The multi wavelet approach can be coupled by a separable form to represent Greens functions of the Poisson and Helmholtz equation [66] which enables efficient Hartree Fock [65] and DFT [64] calculations. These features are implemented in the software package MRChem [60, 61] and this code will be used for all wavelet based calculations in this paper.
@@ III.1 Basis set errors of charge densities
First, the error of the charge density for polarization consistent Gaussian basis sets [45] that are optimized for DFT calculations is analyzed. In Fig. 1, the difference of $\sigma$ between an basis-set-error-free multi wavelet charge density and a charge density obtained with polarization consistent basis sets [45] is shown. The charge density was calculated with the PBE functional for a single neon atom. The “aug” prefix indicates the addition of diffuse functions in the Gaussian basis set. The aug-PC3 and aug-PC4 basis sets agree well with the multi wavelet reference calculation. For comparison, a charge density that was obtained using a correlation consistent basis set (aug-cc-pV5Z) is added in Fig. 1 which gives the same level of accuracy in the charge density. The aug-PC3 contains fewer basis functions than the aug-cc-pV5z basis set, the aug-PC4 more. In Fig. 2 the basis set error of charge densities obtained with a correlation consistent basis [46] sets is shown. Also the aug-cc-pV5Z basis set matches the multi wavelet result almost perfectly. It is therefore recommended to use this basis set whenever highly accurate charge densities are necessary with wave function based methods. The aug-PC3 and aug-PC4 offer the same level with a similar number of basis functions, the aug-PC3 is even a bit smaller than the aug-cc-pV5Z and is therefore an interesting alternative to the correlation consistent basis set when DFT is used. In Fig. 3 it is shown that the aug-cc-pV5Z basis set also produces highly accurate charge densities for coupled cluster calculations with single and double excitations. The error of charge densities with the cc-pVNZ and the PC basis is analyzed in Fig. 4 for a methanol molecule. The reference charge density was obtained using a multi wavelet DFT calculation with the PBE functional. As a measure of error, in this case the absolute value of the difference in Hirshfeld charges is used. Here, the addition of diffuse Gaussian basis functions greatly improves the accuracy and once again, the aug-cc-pV5Z basis is able to produce highly accurate charge densities. The aug-PC3 and aug-PC4 basis sets also produce highly converged Hirshfeld charges, the aug-PC4 appears to be the most accurate basis set in this comparison. However, the calculation using the aug-PC4 basis set was difficult to converge since the condition number of the overlap matrix from the basis functions was larger than $10^{8}$. The aug-PC3, aug-PC4 and the aug-cc-pV5Z bases offer sufficient precision to accurately compare charge densities obtained with different functionals. Given that DFT results are being compared to CCSD calculations, it is more consistent to use the correlation-consistent aug-cc-pV5Z basis set, which is specifically designed for post-Hartree–Fock methods. Therefore, in the remainder of this paper, the aug-cc-pV5Z basis set will be used for comparing various exchange-correlation functionals.
@@ III.2 Accuracy measures of charge densities
The integrals required to calculate the norms were evaluated numerically with the integration grids of PySCF using the accurate grid level. The spherically symmetric reference charge densities required for the calculation of the Hirshfeld charges were computed using Abinit [68]. In Fig. 6, these measures are shown for all functionals. The functionals tested are: